# Event shape dependence of symmetry plane correlations using the Gaussian estimator in Pb-Pb collisions at the LHC using AMPT

### v1

@@ II.3 Event and track selection
The evaluation of SPCs and PPCs as defined in Eq. (II.4) and Eq. (II.4) respectively involves in the RHS, averaging the quantities in the angular brackets over different events. Now, each event is classified as per its centrality, which is done using its impact parameter that goes into the input during event generation, using the Glauber Monte Carlo scheme [43], which is used by the HIJING algorithm during nuclear configuration initialization. Further to this centrality classification, each event is subdivided into an event shape class based on the value of its spherocity. There are three event shape classes in each centrality class – low-$S_{0}$ (pencil-like) events (20% events with the least spherocity), high-$S_{0}$ (isotropic) events (20% events with the highest spherocity) and a pool of events with all possible spherocity values, known as $S_{0}$-int events. The event shape and centrality classifications are listed in Table 1.
@@
After the events have been classified as per their centrality and spherocity classes, the averages in the numerator and denominator in the RHS of Eq. (II.4) are computed using the events that belong to the respective centrality and event shape classes.
The terms inside the angular brackets of the numerator and denominator in the RHS of Eq. (II.4) are obtained for each event. These, in turn, are evaluated by averaging over charged hadron track multiplets as per Eq. (I). For the purpose of averaging over these tracks, only those charged hadronic tracks in the pseudorapidity range $|\eta|<0.8$ and with transverse momenta $5.00.2<p_{\rm T}<$ 5.0 GeV/c are used. These cuts are in order to be consistent with experiment [23] - lower limit owing to ALICE detector efficiency, an upper limit to remove the contribution of non-flow effects that result from jets (which have been taken to be having $p_{\rm T}>5.0$ GeV/c).
II.4 Participant and symmetry plane correlations
From Eq. (I), we can reckon, for a single event, $v_{n}e^{in\psi_{n}}=\langle e^{in\phi}\rangle$. Now, as Pb-Pb collisions at the LHC are associated with high multiplicity, therefore, from event to event, $Re(v_{n})$ and $Im(v_{n})$ (where $Re(z),~{}Im(z)$ denote the real and imaginary part of a complex number $z$ respectively) are expected to have a double Gaussian distribution in the Argand plane [44]. Using this assumption, the correlation between symmetry planes, which are independent of the effect of flow coefficients, has been derived in Ref. [44]. The Gaussian estimator for the symmetry plane correlations is given by [23]:
$\langle\cos(a_{1}n_{1}\psi_{1}+a_{2}n_{2}\psi_{2}+...+a_{k}n_{k}\psi_{k})\rangle_{\rm GE}$
$=\sqrt{\frac{\pi}{4}}\frac{\langle{v_{n_{1}}^{a_{1}}v_{n_{2}}^{a_{2}}...v_{n_{k}}^{a_{k}}\cos(a_{1}n_{1}\psi_{1}+a_{2}n_{2}\psi_{2}+...+a_{k}n_{k}\psi_{k})}\rangle}{\sqrt{\left\langle{v_{n_{1}}^{2a_{1}}v_{n_{2}}^{2a_{2}}...v_{n_{k}}^{2a_{k}}}\right\rangle}}$ (5)
Here, the angular brackets denote event averaging over a suitable set of events in a particular event class 1. Further, following its definition in [44], and as explained in Eq. (3), the choice of harmonics and their powers is made such that $\sum_{i=1}^{k}a_{i}n_{i}=0$. An example of its use and evaluation using multi-particle correlations can be found in Ref. [23].
In the LHS of this equation, $\epsilon_{n}$ are eccentricities of the initial energy distribution, and $\Phi_{n}$, their corresponding symmetry angles, are called the participant plane angle. The RHS shows the way to compute the combination of $\epsilon_{n}$’s and $\Phi_{n}$’s using $\sum_{i=1}^{k}a_{i}=l$ particles. In the RHS, the angles, $\tilde{\phi}_{i}$’s are the azimuthal angles in the coordinate space, of wounded nucleons. Correspondingly, $r_{i}$, which appear to the power $n_{i}$ in the weight factors, are the radial coordinates in the plane azimuthal to the beam axis of the wounded nucleon whose azimuthal angle $\tilde{\phi}_{i}$ is considered. It is important to note that Eq. (II.4) does not entail the correct definition of $\epsilon_{1}e^{i\Phi_{1}}$[45], as well as it fails to account for the minus signs when it comes to computing just one harmonic order of initial state anisotropies. The absence of these features, however, is not an issue for the study of participant plane correlations, as first-order harmonic has not been considered for participant correlations in this paper, and in addition to that, the choice of harmonics $n_{i}$ would lead to all the negative signs canceling when we account for the rotational invariance of the PPCs under the random orientation of the reaction plane from event to event. In case one wishes to study the participant plane correlations involving the first harmonic, i.e., $n=1$, an explicit $r^{3}$ weight factor must be used [45]. Finally, in analogy to Eq. (II.4), we have the following definition of the Gaussian estimator of the PPCs:
$\langle\cos(a_{1}n_{1}\Phi_{1}+a_{2}n_{2}\Phi_{2}+...+a_{k}n_{k}\Phi_{k})\rangle_{\rm GE}$
$=\sqrt{\frac{\pi}{4}}\frac{\langle{\epsilon_{n_{1}}^{a_{1}}\epsilon_{n_{2}}^{a_{2}}...\epsilon_{n_{k}}^{a_{k}}\cos(a_{1}n_{1}\Phi_{1}+a_{2}n_{2}\Phi_{2}+...+a_{k}n_{k}\Phi_{k})}\rangle}{\sqrt{\left\langle{{\epsilon_{n_{1}}^{2a_{1}}\epsilon_{n_{2}}^{2a_{2}}...\epsilon_{n_{k}}^{2a_{k}}}}\right\rangle}}$ (7)
@@ III Results and Discussion
Firstly, it must be noted that, in systems with high multiplicity, the anisotropic flow coefficients provide the extent to which the distribution of particles in the transverse momentum space looks identical to a regular polygon with $n$ sides, where $n$ is the harmonic order of the flow term being considered. The symmetry plane angle denotes the orientation with respect to a reference, which characterizes the particular harmonic order of anisotropy. This is why the correlation between symmetry plane angles of different harmonic orders of anisotropy will provide independent information that can be obtained from the correlation between anisotropic flow coefficients of the same harmonic orders. One particular reason for this is that the anisotropic flow coefficients typically become smaller with the increase in the order of anisotropy, whereas the symmetry planes need not have any such constraints. Hence, a non-zero value of the SPCs would mean that the symmetry plane angles being considered are correlated among themselves. To understand this with an example, we can consider the combination $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$. Consider that the symmetry planes of harmonic orders $n=6$ and $n=3$ are oriented at angles $\delta_{6}$ and $\delta_{3}$, respectively. Clearly, all possible symmetry planes of order 6 are $\{\delta_{6}+\frac{2k\pi}{6}\}_{k=0}^{5}$. Thus, $6\delta_{6}\in\{6\delta_{6}+2k\pi\}_{k=0}^{5}$. Similarly, $6\delta_{3}\in\{6\delta_{3}+4k\pi\}_{k=0}^{2}$. Hence, $\cos(6(\psi_{6}-\psi_{3}))\in\{\cos(6(\delta_{6}-\delta_{3})+(m+2n)2\pi)\}$ for integers $m,n$. The translation of the argument of cosine function by an integer multiple of $2\pi$ leaves the function invariant, and ultimately we can infer that this correlator shows the correlation between any of the orientations of the $6^{\rm th}$ and $3^{\rm rd}$ order anisotropies in the distribution of the transverse momenta of the final state particles that emerge as the result of a collision. Now, further generalizing, a non-zero value of the GE of the SPCs for a linear combination of symmetry planes $\sum_{i=1}^{k}a_{i}n_{i}\psi_{n_{i}}$, indicates that despite event-by-event fluctuations, the orientations of the harmonics $n_{i}$ to their respective powers of $a_{i}$ are correlated. While for simple choice of harmonics, we can provide direct interpretations in terms of geometry, non-trivial behaviour of correlators with centrality has been observed for different choices of $a_{i}$ while keeping $n_{i}$ same under the constraint that $\sum_{i=1}^{k}a_{i}n_{i}=0$.
Figure 1: Comparison of the centrality dependence of SPCs calculated using various models, including AMPT from the present study, with ALICE measurements in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [46].
@@ III.1 Centrality dependence of SPCs
In this section, before explicitly discussing the SPCs, we begin with the comparison of SPCs calculated using AMPT with experimental measurements. Figure 1 shows the comparison of ALICE measurements with corresponding model calculations. Here, one finds that AMPT provides a good qualitative agreement with experimental results while underestimating the values.
@@
Figure 2 shows the symmetry plane correlations of different harmonics as a function of the centrality of collisions in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. In the left panel, which shows the correlations for two symmetry planes, one finds the strength of the correlation is strongest for $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$. The strength of the correlation becomes weaker for $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$, and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$. Further, it becomes the weakest for $\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$, which is consistent with zero within uncertainties. This decreasing strength of correlation between two symmetry planes is consistent with experimental observations made in Refs. [23, 46]. Interestingly, for collision centrality (0-20)%, one finds that $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}>\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}>\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}>\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$. After, (20-30)centrality, $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ becomes comparable and finally $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ leads the values towards (50-60)centrality. Although, different SPCs show different centrality dependence, $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ shows a non-linear centrality dependence. Further, the observed zero value for $\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$ indicates zero correlation between $\psi_{2}$ and $\psi_{3}$, although $v_{2}$ and $v_{3}$ show anti-correlation [23, 46, 47, 48, 49, 50, 51, 52, 53, 20]. This hints towards the applicability of SPCs to independently probe the formation of a QCD medium in heavy-ion collisions. Similar to SPCs involving two symmetry planes, one finds different magnitudes for SPCs of three different symmetry planes. Here, $\langle\cos(2\psi_{2}+3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ shows the strongest correlation and $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ has the weakest correlation. Interestingly, $\langle\cos(2\psi_{2}+3\psi_{3}-6\psi_{6})\rangle_{\rm GE}$ has a negative sign, unlike other correlators which are all positive or consistent with zero at the very least. Another interesting observation is that the magnitudes of the SPCs are approximately dependent to the number of particle correlations in the final state. For example, $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos(2\psi_{2}+3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ each require only three-particle correlations and possesses the largest values among the shown SPCs. In contrast, $\langle\cos[6(\psi_{2}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ require five and six correlators, respectively, and thus possess a negligibly small value of correlation [23, 46]. One possible explanation is as follows. The fluctuation of flow vectors encoded in the SPCs can be traced back to the initial state fluctuations. The initial state fluctuations can be attributed to the fluctuations due to the finite amount of participants in the collision overlap region. According to the central limit theorem, the average of the random samples converges to a Gaussian distribution with an increase in the number of sampling. For a Gaussian distribution, only the second-order cumulant, namely, the width of the distribution, is non-vanishing. The skewness and kurtosis are small for a distribution close to Gaussian. Here, the order of particle correlations refers to the order of cumulants. Thus, the ordering in magnitudes of SPCs shown in Fig. 2 indicates that the contribution of higher order cumulant (higher order particle correlator) is smaller. On the other hand, the lowest-order cumulants have the leading values. This explains the crossing behaviour between $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$, after (20-30)centrality class, where the number of participating nucleons is smaller. We now move to discuss the geometrical interpretation of each SPC and corresponding PPC.
@@
Figure 3: Centrality dependence of SPCs (upper panel) among $n=2,3,6$ and corresponding PPCs (lower panel) in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. The calculation of SPCs considers the charged hadrons with $5.00.2<p_{\rm T}<$ 5.0 GeV/c, $|\eta|<0.8$.
@@ III.2.1 SPCs of harmonic orders $n=2,3,6$
The symmetry planes corresponding to the harmonics $n=2,3,6$ represent elliptical, triangular, and hexagonal symmetries, respectively. The correlations amongst them have been shown in Figs. 3(a)-(c). Whenever there is a hexagonal symmetry present in the system, we can deduce trivially from the geometry that choosing any two diametrically opposite vertices of a regular hexagon (that is, those vertices that are separated by the largest distance) would also correspond to a local elliptical symmetry about them. Hence, it can be expected that there is a second-order symmetry plane whose alignment is correlated to the sixth-order symmetry plane. This will be an enhanced effect when $v_{6}$ is more dominant than the second-order anisotropy. In such cases, the symmetry planes $\psi_{6}$ and $\psi_{2}$ are more likely to be correlated. Similarly, by choosing any of the three vertices of a regular hexagon that alternate with each other on the hexagon, we obtain an equilateral triangle. So, a plane about which the transverse distribution has a hexagonal symmetry is also going to be correlated to a plane corresponding to triangular symmetry. Thus, in addition to a strong correlation in $\psi_{6}$ and $\psi_{2}$, a strong correlation between $\psi_{6}$ and $\psi_{3}$ can also be expected for events with higher values of $v_{6}$ than that of $v_{3}$. Again, along the same line of deduction, a higher $v_{6}$ than both $v_{2}$ and $v_{3}$ would also correspond to a higher correlation between $\psi_{2}$ and $\psi_{3}$. Another interesting feature that follows from the presence of hexagonal symmetry is that there are three distinct diametrically opposite points and two distinct equilateral triangles that can be constructed using the six vertices of a regular hexagon. And more generally, given the geometrical fact that for natural numbers $m$ and $n$, we can find $m$ diameters in $2m$ sided regular polygons and $n$ equilateral triangles using the vertices of a $3n$ sided regular polygon only, it follows that in systems where anisotropic flows of order $6k$ dominate, for a natural number $k$, we can expect that the SPCs $\langle\cos[6k(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos[6k(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ would have non-vanishing magnitudes with varying centrality and spherocity. The converse will also hold true, i.e., if the SPCs $\langle\cos[6k(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos[6k(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ are comparable and non-vanishing, then $v_{6k}$ is dominant in the system for some natural number $k$, since there exist no other regular polygons, in which, few vertices can be chosen ensuring that the straight line connecting them is either a diameter or that they form an equilateral triangle. For the case when $k=1$, the above discussion holds true for hexagonal symmetry in the final state transverse momentum distribution. As can be seen, however, in Fig. 3(a), for the most central event classes, there is an overlap between the values of the SPCs in the different event shape classes. This is indicative of the fact that this correlator takes on its values due to event-by-event fluctuations in the Pb-Pb collision system rather than due to any dominant sixth-order anisotropy. More particularly, comparing Fig. 5(a) and Fig. 3(b), we can observe that although the SPC $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ has a qualitatively similar centrality dependence as that of $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$ for the mid-central classes, it has different behaviour in the most central classes. From the inference previously drawn from $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$, it can be understood that the values $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ takes is not due to an underlying anisotropy with hexagonal symmetry, but that particle emission in Pb-Pb collisions is largely dominated by $\psi_{2}$, as all other anisotropic flow terms only have a weak contribution.
@@
However, if second and third-order anisotropies dominate in the system, then no such trivial link can be deduced. More specifically, if $v_{2}$ or $v_{3}$ dominate in comparison to $v_{6}$, then there are a few $6^{\rm th}$ order symmetry planes close to which the $2^{\rm nd}$ and $3^{\rm rd}$ order symmetry planes can lie in the transverse momentum space, which is the case for of Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. Although from the above discussion, the magnitude of the anisotropic flow coefficients seems to be playing an important role in the value that the SPCs take on, it is important to note that the anisotropic flow coefficient magnitudes have only been used to represent the symmetry of the anisotropic flow, and their contribution to SPCs has been removed by suitably choosing the harmonics and their respective powers in Eq. (II.4).

### v2

@@ II.4 Event and track selection
The evaluation of SPCs and PPCs as defined in Eq. (5) and Eq. (7), respectively involve in the RHS, averaging the quantities in the angular brackets over different events. Now, each event is classified as per its centrality, which is done using its impact parameter that goes into the input during event generation. Glauber Monte Carlo scheme [51], which is used by the HIJING algorithm during nuclear configuration initialization, is also used for centrality classification. Further to this centrality classification, each event is subdivided into an event shape class based on the value of its spherocity. There are three event shape classes in each centrality class – low-$S_{0}$ (pencil-like) events (20% events with the least spherocity), high-$S_{0}$ (isotropic) events (20% events with the highest spherocity) and a pool of events with all possible spherocity values, known as $S_{0}$-int events. The event shape and centrality classifications are listed in Table 1.
@@
After the events have been classified into their centrality and spherocity classes, the event averages in the numerator and denominator in the RHS of Eq. (5) are computed one by one for these different event classes.
The terms inside the angular brackets of the numerator and denominator in the RHS of Eq. (5) are obtained for each event. These, in turn, are evaluated by averaging over the charged hadron track multiplets as per Eq. (I). For the purpose of averaging over these tracks, only those charged hadronic tracks in the pseudorapidity range $|\eta|<0.8$ and with transverse momenta $0.2<p_{\rm T}<5.0$ GeV/c are used. These cuts are in order to be consistent with experiment [23] - lower limit owing to ALICE detector efficiency, and upper limit to remove the contribution of non-flow effects that result from jets (which is taken to be $p_{\rm T}>5.0$ GeV/c).
@@ III Results and Discussion
Firstly, it must be noted that, in systems with high multiplicity, the anisotropic flow coefficients of harmonic order $n$ quantify the extent to which the distribution of particles in the transverse momentum space looks identical to a regular polygon with $n$ sides. The anisotropic flow coefficients typically become smaller with the increase in the order of anisotropy. So, the studies of correlations amongst the anisotropic flow coefficients provide us with insight into the development of flow in the collision system.
The symmetry plane angle of a harmonic order $n$, on the other hand, denotes the orientation with respect to a reference, which characterizes the particular harmonic order of anisotropy. So, the symmetry planes need not vanish, as we go on to study the higher harmonics. Hence, a non-zero value of the SPCs would mean that the symmetry plane angles being considered are correlated among themselves, indicating in turn that the corresponding harmonic orders of flow are correlated. This is why the correlation between symmetry plane angles of different harmonic orders of anisotropy will provide independent information, than what can be obtained from the correlation between anisotropic flow coefficients of the same harmonic orders.
To understand this with an example, we can consider the combination $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$. Consider that the symmetry planes of harmonic orders $n=6$ and $n=3$ are oriented at angles $\delta_{6}$ and $\delta_{3}$, respectively. Clearly, all possible symmetry planes of order 6 are $\{\delta_{6}+\frac{2k\pi}{6}\}_{k=0}^{5}$. Thus, $6\delta_{6}\in\{6\delta_{6}+2k\pi\}_{k=0}^{5}$. Similarly, $6\delta_{3}\in\{6\delta_{3}+4k\pi\}_{k=0}^{2}$. Hence, $\cos(6(\psi_{6}-\psi_{3}))\in\{\cos(6(\delta_{6}-\delta_{3})+(m+2n)2\pi)\}$ for integers $m,n$. The translation of the argument of the cosine function by an integer multiple of $2\pi$ leaves it’s value invariant, and ultimately we can infer that this correlator gives the correlation between any of the orientations of the $6^{\rm th}$ and $3^{\rm rd}$ order symmetry planes in the distribution of the transverse momenta of the final state particles that emerge from of a collision. Now, further generalizing, a non-zero value of the GE of the SPCs for a linear combination of symmetry planes $\sum_{i=1}^{k}a_{i}n_{i}\psi_{n_{i}}$, indicates that, despite event-by-event fluctuations, the orientations of the harmonics $n_{i}$ to their respective powers of $a_{i}$ are correlated. While for simple choice of harmonics, we can provide direct interpretations in terms of geometry of the event in momentum space, non-trivial behaviour of correlators with centrality has been observed for different choices of $a_{i}$, keeping $n_{i}$ the same under the constraint that $\sum_{i=1}^{k}a_{i}n_{i}=0$.
Figure 1: Comparison of the centrality dependence of SPCs calculated using various models, including AMPT from the present study, with ALICE measurements in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [58]. Corresponding measurements from TRENTO, EKRT+hydrodynamics [52, 53], TRENTO+iEBE-VISHNU [54, 55, 56, 57] are taken from Ref. [58].
@@ III.1 Centrality dependence of SPCs
In this section, before explicitly discussing the SPCs, we begin with the comparison of the SPCs of the present study, calculated using AMPT, with experimental measurements from ALICE and other hydrodynamical models, shown in Fig. 1. The measurements from TRENTO, EKRT+hydrodynamics [52, 53], TRENTO+iEBE-VISHNU [54, 55, 56, 57] are taken from Ref. [58]. Here, one finds that AMPT provides a good qualitative agreement with the experimental results while underestimating the values, while the hydrodynamical models overestimate the experimental findings. This hints that the SPCs have strong constraints on the model details.
@@
Figure 2 shows the symmetry plane correlations of different harmonics as a function of the centrality of collisions in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. In the left panel, which shows the correlations for two symmetry planes, one finds that the strength of the correlation is strongest for $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$. The correlation becomes weaker for $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$. The correlation strength is weakest in the central collisions due to the isotropic geometry of the collision overlap region. The values of these SPCs start to rise towards the mid-central collisions, which possess contributions from initial collision geometry. Moreover, due to a smaller number of participants, the values of these SPCs start to decrease towards the peripheral collisions. This decreasing strength of correlation between these symmetry planes is consistent with experimental observations made in Refs. [23, 58].
Interestingly, for collision centrality (0-20)%, one finds that $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}>\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}>\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}>\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$. After, (20-30)% centrality, $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ become comparable and finally $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ leads the values towards (50-60)% centrality. Although, different SPCs show different centrality dependence, $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ shows a non-linear centrality dependence. Further, the observed zero value for $\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$ indicates zero correlation between $\psi_{2}$ and $\psi_{3}$, although $v_{2}$ and $v_{3}$ show anti-correlation [23, 58, 59, 60, 61, 62, 63, 64, 52, 20]. This hints towards the applicability of SPCs to independently probe the formation of a QCD medium in heavy-ion collisions.
Similar to SPCs involving two symmetry planes, one finds different magnitudes for SPCs of three different symmetry planes. Here, $\langle\cos(2\psi_{2}+3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ shows the strongest correlation and $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ has the weakest correlation. Interestingly, $\langle\cos(2\psi_{2}+3\psi_{3}-6\psi_{6})\rangle_{\rm GE}$ has a negative sign, unlike other correlators, which are all positive or consistent with zero at the very least.
@@
Another interesting observation is that the magnitudes of the SPCs are approximately dependent on the number of particle correlations in the final state. For example, $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos(2\psi_{2}+3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ each require only three-particle correlations and possess the largest values among the shown SPCs (Fig. 2). In contrast, $\langle\cos[6(\psi_{2}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ require five and six particle correlations, respectively, and thus possess a negligibly small value [23, 58]. One possible explanation is as follows. The fluctuation of flow vectors encoded in the SPCs can be traced back to the initial state fluctuations. The initial state fluctuations can be attributed to the fluctuations due to the finite number of participants in the collision overlap region. According to the central limit theorem, the average of the random samples converges to a Gaussian distribution with an increase in the number of sampling. For a Gaussian distribution, only the second-order cumulant, namely, the width of the distribution, is non-vanishing. The skewness and kurtosis are small for a distribution close to Gaussian. Here, the order of particle correlations refers to the order of cumulants. Thus, the ordering in magnitudes of SPCs shown in Fig. 2 indicates that the contribution of the higher order cumulant (higher order particle correlator) is smaller. On the other hand, the lowest-order cumulants have the leading values. This explains the crossing behaviour between $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$, after (20-30)% centrality class, where the number of participating nucleons is smaller.
We now move to discuss the geometrical interpretation of each SPC and its corresponding PPC in the following section.
@@
Figure 3: Centrality dependence of SPCs (shown in (a), (b), and (c)) among $n=2,3,6$ and corresponding PPCs (shown in (d), (e), and (f)) in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. The calculation of SPCs considers the charged hadrons with $0.2<p_{\rm T}<5.0$ GeV/c, $|\eta|<0.8$.
@@ III.2.1 SPCs of harmonic orders $n=2,3,6$
The symmetry planes corresponding to the harmonics $n=2,3,6$ represent elliptical, triangular, and hexagonal symmetries, respectively. The correlations amongst them have been shown in Figs. 3(a)-(c). Whenever there is a hexagonal symmetry present in the system, we can deduce trivially from the geometry that choosing any two diametrically opposite vertices of a regular hexagon (that is, those vertices that are separated by the largest distance) would also correspond to a local elliptical symmetry about them. Hence, it can be expected that there is a second-order symmetry plane whose alignment is correlated to the sixth-order symmetry plane. This gives rise to the presence of elliptic symmetry, which contributes nonlinearly in the cumulant expansion of sixth-order harmonics [50, 65, 66, 58]. Similarly, by choosing any of the three vertices of a regular hexagon that alternate with each other on the hexagon, we obtain an equilateral triangle. So, a plane about which the transverse distribution has a hexagonal symmetry is also going to be correlated to a plane corresponding to triangular symmetry, i.e., the presence of non-linear triangular symmetry in the cumulant expansion of hexagonal harmonics. Thus, in addition to a strong correlation in $\psi_{6}$ and $\psi_{2}$, a strong correlation between $\psi_{6}$ and $\psi_{3}$ can also be expected. Again, along the same line of deduction, with a rare occurrence of higher $v_{6}$ than both $v_{2}$ and $v_{3}$ would correspond to a higher correlation between $\psi_{2}$ and $\psi_{3}$, which is not usually expected otherwise.
@@
Another interesting feature that follows from the presence of hexagonal symmetry is that there are three distinct diametrically opposite points and two distinct equilateral triangles that can be constructed using the six vertices of a regular hexagon. And more generally, given the geometrical fact that for natural numbers $m$ and $n$, we can find $m$ diameters in $2m$ sided regular polygons and $n$ equilateral triangles using the vertices of a $3n$ sided regular polygon only, it follows that in system with finite $6k$ order harmonics, for a natural number $k$, we can expect that the SPCs $\langle\cos[6k(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos[6k(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ would have non-vanishing magnitudes with varying centrality and spherocity. The converse will also hold true, i.e., if the SPCs $\langle\cos[6k(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos[6k(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$ are comparable and non-vanishing, then $v_{6k}$ is non-zero in the system for some natural number $k$. This is because there exist no other regular polygons, in which a few vertices can be chosen, ensuring that the straight line connecting them is either a diameter or that they form an equilateral triangle. For the case when $k=1$, the above discussion holds true for hexagonal symmetry in the final state transverse momentum distribution.
As shown in Fig. 3(a), for the most central event classes, there is an overlap in $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle$ among different event shape classes. This is indicative of the fact that this correlator takes on its values due to event-by-event fluctuations in the central Pb-Pb collisions, which is overall a reflection of corresponding PPCs shown in Fig. 3 (d). More particularly, comparing Fig. 5(a) and Fig. 3(b), we can observe that although the SPC $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ has a qualitatively similar centrality dependence as that of $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$ for the non-central classes, it has a different behaviour in the most central classes. From the inference previously drawn from $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$, it can be understood that the values $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ for the non-central classes, is not due to an underlying anisotropy with hexagonal symmetry, but that particle emission in Pb-Pb collisions is largely dominated by $\psi_{2}$, and all other harmonic terms only have a weak contribution.
@@
However, if second and third-order anisotropies dominate in the system, no such trivial link can be deduced. More specifically, if $v_{2}$ or $v_{3}$ dominate in comparison to $v_{6}$, then there are a few $6^{\rm th}$ order symmetry planes close to which the $2^{\rm nd}$ and $3^{\rm rd}$ order symmetry planes can lie in the transverse momentum space. This is precisely the case for Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. Although from the above discussion, the magnitude of the anisotropic flow coefficients seems to be playing an important role in the value that the SPCs take on, it is important to note that their magnitudes have only been used to characterize the dominant symmetries of the anisotropic flow, and their contribution to SPCs has been removed by suitably choosing the harmonics and their respective powers in Eq. (5).
@@
To make it clearer, a comparison has been shown in Fig. 4 between the SPC $\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$, which has been obtained using all events in a centrality class, to that obtained using events with higher values of $v_{6}$. We can observe no positive shift in the values, owing to the geometrical fact that the particles are not ejected with transverse momenta distributed in a hexagonal symmetry.